If an object is thrown vertically upwards with speed 49 ms-1. How long does it take to
complete upward journey? What maximum height does it achieve?

Grade:9

2 Answers

Aarti Gupta
9 years ago
Given-
Initial velocity,u = 49m/s
Acceleration due to gravity,g = -9.8m/s2 (as object is thrown upwards against gravity)
Final velocity,v = 0 as upon attainig maxm. height final velocity gets = 0
Time taken,t = ?
Maxm. height,h = ?
Solution-
From first euqation of motion,
v = u + gt
0 = 49 – 9.8t
t = 49/9.8 = 5s
t = 5s
From third equation of motion,we have,
v2 = u2 + 2gh
0 = 492 – 2*9.8*h
h = 122.5m
Thus,time taken to complete upward journey,t = 5s and maxm. height attained by object,h= 122.5m
MRUDUL APPARI
9 years ago
thank you